OpenStudy (barrelracing):

Given the function f(x) = x2 and k = –3, which of the following represents a vertical shift? a) f(x) + k b) kf(x) c) f(x + k) d) f(kx)

OpenStudy (solomonzelman):

you don't need to know the value of k to answer this question. you know that when you have y=f(x) you get certain y-coordinates, but when you add k (like this: y=f(x)+k) then you know that for each x-coordinate you get an output +k greater than in a case of y=f(x), and that means that f(x)+k is a shift up)

OpenStudy (solomonzelman):

(for k<0. tho' , like in your case where x=-3, it is a shift down)

OpenStudy (solomonzelman):

For any parent function y=f(x), the rule is: ------------------- VERTICAL SHIFTS y=f(x)+c shift c units up y=f(x)-c shift c units down ------------------- SIDE SHIFTS y=f(x-c) shift c units to the right y=f(x+c) shift c units to the left -------------------